OpenStudy (anonymous):

the track at cassie school is 2 3/5 miles. cassie ran around the track 2 3/4 times. how many miles did she run the track at cassie school is 2 3/5 miles. cassie ran around the track 2 3/4 times. how many miles did she run @Mathematics

OpenStudy (anonymous):

2 3/5 * 2 3/4 = 13/5 * 11/4 = 143/20 = 7 3/20 miles

OpenStudy (anonymous):

so is that the answer

OpenStudy (anonymous):

thanks i will give u medal